OpenStudy (anonymous):

Find the points on the graph of y=4-x^2 that are closest to the point (0,2)

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

I think you'll have to use the distance formula. let x2 be x, x1 = 0, y2 = 4-x^2, y1 = 2 \[d ^{2}=(x-0)^{2}+((4-x ^{2})-2)^{2}\]

OpenStudy (anonymous):

k

OpenStudy (agent0smith):

And then you want to minimize that, so differentiate and set it to zero

OpenStudy (anonymous):

\[d=\sqrt{(x-0)^2 + (y-2)^2}\]

OpenStudy (anonymous):

\[d=\sqrt{x^2+(4-x^2-2)^2}\]

OpenStudy (agent0smith):

Just leave it as d^2, since if you minimize d^2, you also minimize d. Also the square root will disappear later anyway, even if you did use it.

OpenStudy (anonymous):

\[d=\sqrt{x^2+(2-x^2)^2}\]

OpenStudy (anonymous):

\[d'=\frac{ 1 }{ 2 } \times (x^2+(2-x^2)^2)^\frac{ -1 }{ 2 } \times 2x +2(2-x^2)(-2x)\]

OpenStudy (anonymous):

\[\frac{ 2x-4x(2-x^2) }{ 2\sqrt{x^2+(2-x^2)^2)} } = 0\]

OpenStudy (anonymous):